# A Framework for Lorentz-Dirac Dynamics and Post-Newtonian Interaction of Radiating Point Charges

### v4

@@ Abstract
We examine classical radiation reaction by combining the covariant Lorentz–Dirac formulation, its Landau–Lifshitz (LL) order reduction, and a post-Newtonian (PN) Hamiltonian treatment of interacting and radiating charges. After reviewing the LL reduction and its removal of runaway and preacceleration behavior, we verify energy balance in several relativistic single-particle scenarios by demonstrating agreement between the LL Larmor power and the loss of mechanical energy. We then construct an N-body framework based on the conservative Darwin Hamiltonian supplemented with the leading 1.5PN radiation–reaction term. Numerical simulations of charge-neutral binary systems of both symmetric and asymmetric mass configurations show orbital decay, circularization, and monotonic Hamiltonian decrease consistent with dipole radiative losses. The resulting framework provides a simple analogue of gravitational PN radiation reaction and a tractable system for studying dissipative and potentially chaotic electromagnetic dynamics.
@@ I Introduction
The motion of an accelerating charge is influenced not only by externally applied electromagnetic fields, but also by the particle’s own field. The resulting radiation–reaction (self–force) problem has a century–long history and remains central to the modeling of high–energy astrophysical systems in which radiation losses shape the dynamics of relativistic plasmas. In pulsar and magnetar magnetospheres, black–hole magnetospheres, and gamma–ray burst outflows, charged particles are routinely accelerated to ultra–relativistic energies and radiate copiously through curvature, synchrotron, and inverse–Compton processes, thereby regulating both particle distributions and observed spectra [1, 2, 3, 4]. A systematic classical description of radiation reaction is also valuable as an analogue system for gravitational self–force and waveform modeling in compact binaries, where back–reaction of emitted gravitational waves drives inspiral and determines the phase evolution of signals observed by ground– and space–based detectors [5, 6].
In relativistic electrodynamics the Lorentz–Dirac (LD) equation supplements the Lorentz force with a covariant self–force term linear in the proper–time derivative of the acceleration, capturing the leading effect of radiation back–reaction on the worldline of a point charge [7, 8]. The LD equation admits complementary derivations: a decomposition of the electromagnetic field into half–retarded minus half–advanced parts that isolates its radiative component [10, 9], and Dirac’s world–tube momentum–balance argument based on the electromagnetic stress–energy tensor [7, 8]. Both approaches emphasize the split between bound and radiative degrees of freedom and lead to the same covariant structure. However, the LD equation is third order in time and exhibits well–known pathologies such as runaway solutions and preacceleration, reflecting the need to prescribe initial acceleration data in addition to position and velocity [11, 12, 13]. These issues are resolved by the reduction–of–order prescription of Landau and Lifshitz, which treats the self–force as a perturbation and replaces the derivative of the acceleration by the derivative of the external Lorentz force. The resulting Landau–Lifshitz (LL) equation is second order, causal, and has been justified rigorously in modern self–force treatments of extended charged bodies in the point–particle limit [14, 15, 16].
Although historically motivated by internal consistency of classical electrodynamics, radiation-reaction dynamics has acquired renewed practical importance. In high–energy astrophysics, self–forces govern the damping of particle orbits in pulsar winds, regulate curvature and synchrotron emission in strongly magnetized neutron stars, and influence nonthermal spectra in relativistic jets and gamma–ray bursts [1, 2, 3, 4]. In parallel, ultra-intense laser facilities now access electromagnetic fields comparable to those near compact objects, enabling laboratory studies of radiation reaction in the collision of relativistic electron beams with petawatt–class laser pulses [17]. Recent experiments have reported direct signatures of radiation–reaction–induced energy loss and high–energy photon emission, in good agreement with classical and semiclassical models based on LL–type dynamics [18, 19]. In this work we view such laser applications as a valuable secondary arena, while placing primary emphasis on the role of electromagnetic self–forces as a useful analogue for gravitational radiation reaction in high energy astrophysical scenarios.
The gravitational–wave community has developed a sophisticated post–Newtonian (PN), self–force, and effective–one–body (EOB) formalism to describe the conservative and dissipative dynamics of compact binaries. In PN and ADM Hamiltonian formulations, the conservative sector is encoded in a canonical Hamiltonian known through at least 3PN–4PN order [20, 21, 22], while dissipative radiation–reaction terms first enter at 2.5PN and 3.5PN order [23, 24, 25, 26]. The EOB formalism reorganizes this information into a resummed Hamiltonian and a nonconservative force added to the momentum equation, calibrated to PN fluxes and numerical–relativity waveforms [27, 28, 29, 30, 31, 32]. Effective field theory (EFT) approaches further show that such radiation–reaction forces arise naturally from a doubled–field nonconservative action principle [33, 34]. An electromagnetic analogue that mirrors the PN/EOB “Hamiltonian plus radiation–reaction” framework can provide useful intuition and a testbed for ideas in gravitational–wave theory [5, 6, 16]. While radiation reaction, Landau–Lifshitz dynamics, and post-Newtonian expansions have each been studied extensively, an explicit Hamiltonian-plus-radiation-reaction framework for interacting charged particles in electrodynamics—directly analogous to PN/EOB treatments of gravitational binaries—has remained largely unexplored.
In this paper we construct such an analogue by combining the covariant LD/LL description with a PN Hamiltonian treatment of interacting charges. We first revisit the LD and LL equations, giving concise derivations based on the bound/radiative field split and on Dirac’s world–tube momentum balance, and we review the appearance of runaway and preaccelerated solutions together with their removal by reduction of order [7, 10, 9, 8, 14, 15]. We then implement the LL equation in several relativistic single–particle scenarios (cyclotron motion and three representative external–field configurations) and verify energy balance by comparing the change in relativistic kinetic energy, the work done by external fields, and the covariant Larmor power. Motivated by PN and EOB literature, we subsequently construct an NN–body framework in which the conservative dynamics are generated by the Darwin Hamiltonian at 1PN order, while the leading dipole radiation–reaction effects enter as a nonconservative 1.5PN force in the momentum equation, derived via order reduction in close analogy with gravitational PN radiation reaction [23, 24, 26, 33]. Numerical simulations of charge neutral binary systems of equal–mass and asymmetric mass systems then demonstrate orbital decay, circularization, trajectory deformation, and monotonic Hamiltonian decrease consistent with dipole Larmor losses. Taken together, these results provide a useful electromagnetic analogue of PN gravitational radiation reaction and a tractable system for exploring dissipative and potentially chaotic dynamics relevant to high–energy astrophysical plasmas, gravitational–wave source modeling, and strong–field laser–plasma experiments.
@@
The structure of this paper is as follows. In Sec. I we introduce the radiation–reaction problem in classical electrodynamics, Section II reviews the Lorentz–Dirac equation and its formulation in terms of bound and radiative field components, while Sec. III presents the Landau–Lifshitz reduction–of–order procedure that yields a causal, second–order evolution equation. In Sec. IV we validate this reduced dynamics through numerical simulations of single–particle motion in static, time–dependent, and multidimensional electromagnetic fields. Section V introduces a post–Newtonian truncation of the Lorentz–Dirac force suitable for multi–particle systems, and Sec. VI develops the combined Darwin (1PN) conservative Hamiltonian and the leading 1.51.5PN dipole radiation–reaction term, together with the corresponding Hamilton equations. In Sec. VII, we present numerical simulations of binary charge-neutral systems subject to both long–range Coulomb interactions and dissipative radiation reaction, illustrating orbital decay, circularization, deformation, and the onset of non–integrable dynamics. We summarize the main results and outline future directions in Sec. VIII. Several technical details are collected in the Appendices: App. A reviews the runaway solutions of the Lorentz–Dirac equation, App. B summarizes electromagnetic field–tensor and metric conventions, App. C provides explicit expressions for the Darwin 1PN corrections, and App. D gives a more detailed discussion of the four–acceleration and the covariant Larmor power.
II Radiation Reaction and the Modified Lorentz–Dirac Equation
The Lorentz–Dirac equation provides the classical expression for the radiation-reaction force on a point charge and admits two standard derivations. The first, following Landau and Lifshitz [10, 8], identifies the self-field as the half-retarded minus half-advanced solution, isolating the radiative component of the electromagnetic field. The second, due to Dirac [7] and developed further in [9], derives the same equation from energy–momentum conservation in a world-tube surrounding the charge. Both approaches lead to the covariant Lorentz–Dirac equation,
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta},$ (1)
which contains third-order derivatives of the worldline. Here $F^{\alpha}_{\mathrm{ext}}$ denotes an externally applied (non-self) force acting on the particle. The term in parentheses, $\left(\delta^{\alpha}_{\beta}+\dfrac{u^{\alpha}u_{\beta}}{c^{2}}\right)$, acts as a projection operator orthogonal to the four–velocity. It ensures that the self–force satisfies the constraint $u_{\alpha}a^{\alpha}=0$, so that the total four–acceleration remains orthogonal to $u^{\alpha}$, consistent with the normalization condition $u_{\alpha}u^{\alpha}=-c^{2}$. As is well known, Eq. (1) admits unphysical runaway and preaccelerated solutions. To remedy this, Landau and Lifshitz proposed a reduction-of-order procedure in which $\dot{a}^{\alpha}$ is replaced by the proper-time variation of the external force, yielding the physically meaningful equation of motion
@@ II Radiation Reaction and the Modified Lorentz–Dirac Equation
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}mc^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)F^{\beta}_{\mathrm{ext},\gamma}u^{\gamma},$ (2)
free of runaway solutions and preacceleration, and now regarded as the correct classical description of radiation reaction [13, 12, 11]. Before delving into the dynamics prescribed by the above equation, let us have a brief recap on the Dirac-Lorentz equation based on the Landau-Lifshitz approach and next, the Dirac approach.
II.1 Landau Lifshitz approach
The electromagnetic field generated by an accelerating point charge acts back upon the charge itself, producing what is known as the radiation-reaction or self-force. The classical and systematic approach to deriving this effect, following Landau and Lifshitz [10, 8] begins by decomposing the electromagnetic potential into retarded and advanced solutions of the wave equation,
$\square A^{\alpha}=-4\pi j^{\alpha},$ (3)
where $j^{\alpha}$ is the current of the point charge moving along a worldline $z^{\alpha}(\tau)$. Let $A^{\alpha}_{\mathrm{ret}}$ denote the physically causal field and $A^{\alpha}_{\mathrm{adv}}$ its time-reversed counterpart. Their symmetric and antisymmetric combinations,
$A^{\alpha}_{\mathrm{sym}}=\frac{1}{2}\left(A^{\alpha}_{\mathrm{ret}}+A^{\alpha}_{\mathrm{adv}}\right),\qquad A^{\alpha}_{\mathrm{rr}}=\frac{1}{2}\left(A^{\alpha}_{\mathrm{ret}}-A^{\alpha}_{\mathrm{adv}}\right),$ (4)
have distinct physical interpretations [7, 9]. The symmetric part represents the Coulomb-like near field bound to the particle, while the antisymmetric combination
$A^{\alpha}_{\mathrm{rr}}$ (5)
contains the radiative information responsible for energy loss and self-force. The corresponding electromagnetic field tensor is
$F^{\alpha\beta}_{\mathrm{rr}}=\partial^{\alpha}A^{\beta}_{\mathrm{rr}}-\partial^{\beta}A^{\alpha}_{\mathrm{rr}}.$ (6)
In the momentarily comoving Lorentz frame (MCLF), where the charged particle is instantaneously at rest, one finds the leading contribution to the radiation-reaction field,
$\bm{E}_{\mathrm{rr}}=\frac{2}{3}\,\frac{q}{4\pi\varepsilon_{0}c^{3}}\,\dot{\bm{a}},\qquad\bm{B}_{\mathrm{rr}}=\bm{0},$ (7)
where $\bm{a}$ is the instantaneous acceleration and $\dot{\bm{a}}$ its proper-time derivative [11, 12]. Substitution into the Lorentz force law gives the familiar nonrelativistic expression,
$\bm{F}_{\mathrm{rr}}=\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\dot{\bm{a}}.$ (8)
The covariant generalization of (8), uniquely consistent with $u_{\alpha}a^{\alpha}=0$, is
$F^{\alpha}_{\mathrm{rr}}=\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta}.$ (9)
Including an external force $F^{\alpha}_{\mathrm{ext}}$, the full equation of motion becomes
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta},$ (10)
which is the Lorentz–Dirac equation. As is well known, Eq. (10) involves third derivatives of the position and admits unphysical runaway and pre-accelerated solutions. A consistent physical formulation requires the reduction-of-order procedure of Landau and Lifshitz [10, 13], as we will discuss in Sec. III.
II.2 Dirac’s Approach
A second and conceptually distinct route to the Lorentz–Dirac equation is due to Dirac [7], based on the conservation of total energy–momentum for the combined system of particle and electromagnetic field. In this formulation, one analyzes the flux of electromagnetic stress–energy through a narrow world tube enclosing the particle’s worldline, and equates this to the change in the particle’s mechanical momentum. Modern geometric expositions of this method are given in [9, 8].
II.2.1 World-Tube Construction and Momentum Balance
Let $T^{\alpha\beta}_{\mathrm{em}}$ denote the electromagnetic stress–energy tensor, and surround the particle’s worldline $z^{\alpha}(\tau)$ with a world tube $\Sigma$ of small radius $r$. The flow of electromagnetic four-momentum across this surface is
$\frac{dP^{\alpha}_{\mathrm{em}}}{d\tau}=\int_{\Sigma}T^{\alpha\beta}_{\mathrm{em}}\,d\Sigma_{\beta},$ (11)
where $d\Sigma^{\beta}$ is the outward-directed hypersurface element. Because $T^{\alpha\beta}_{\mathrm{em}}$ is divergenceless in the vacuum exterior of the worldline, the result is independent of the precise choice of tube.
The retarded electromagnetic field naturally decomposes into a near-field bound component and a radiative component [9]. Evaluating (11) yields
$\frac{dP^{\alpha}_{\mathrm{em}}}{d\tau}=\frac{q^{2}}{8\pi\varepsilon_{0}rc^{2}}\,a^{\alpha}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\frac{a^{2}}{c^{2}}\,u^{\alpha},$ (12)
where $u^{\alpha}$ and $a^{\alpha}=du^{\alpha}/d\tau$ are the four-velocity and four-acceleration of the particle, and $a^{2}=a_{\beta}a^{\beta}$. The first term diverges as $r\to 0$ and is interpreted as contributing to the particle’s inertial mass (electromagnetic mass). The second term, finite and directed along $u^{\alpha}$, corresponds to the radiated energy–momentum.
II.2.2 Mechanical Momentum and Mass Renormalization
The total momentum of the particle-plus-field system must be conserved in the absence of exterbnal forces:
$\frac{d}{d\tau}\!\left(P^{\alpha}_{\mathrm{mech}}+P^{\alpha}_{\mathrm{em}}\right)=F^{\alpha}_{\mathrm{ext}}.$ (13)
The divergent term in (12) is absorbed via mass renormalization,
$m=m_{0}+\frac{q^{2}}{8\pi\varepsilon_{0}rc^{2}}\,$ (14)
and consistency with $u_{\alpha}u^{\alpha}=-1$ requires the mechanical momentum to take the form
$P^{\alpha}_{\mathrm{mech}}=m_{0}u^{\alpha}-\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}\,c^{3}}\,a^{\alpha}.$ (15)
The minus sign in $P^{\alpha}_{\rm mech}$ is fixed by the requirement $u_{\alpha}u^{\alpha}=-c^{2}$: because $\dot{P}^{\alpha}_{\rm em}\propto\frac{2}{3}\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\frac{a^{2}}{c^{2}}u^{\alpha}$, total momentum conservation demands $P_{\rm mech}\!\cdot a=-\frac{2}{3}\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}a^{2}$ [8].
Substituting (12) and (15) into (13) produces the covariant Lorentz–Dirac equation,
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\left(\dot{a}^{\alpha}-\frac{a^{2}u^{\alpha}}{c^{2}}\right),$ (16)
where $F^{\alpha}_{\mathrm{ext}}=qF^{\alpha\beta}_{\mathrm{ext}}u_{\beta}$ is the external Lorentz force. Because the four-acceleration is orthogonal to the four-velocity,
$u_{\alpha}a^{\alpha}=0,$ (17)
differentiating this orthogonality relation with respect to proper time $\tau$ yields
$\frac{d}{d\tau}(u_{\alpha}a^{\alpha})=\dot{a}_{\alpha}u^{\alpha}+a_{\alpha}a^{\alpha}=0.$ (18)
Therefore,

### v5

# Post-Newtonian Dynamics of Radiating Charges: Canonical Formulation and Binary Inspiral Laws

@@ Abstract
We construct an explicit electromagnetic analogue of the post-Newtonian Hamiltonian framework widely used in gravitational-wave physics. Starting from the Lorentz–Dirac equation and implementing Landau–Lifshitz order reduction, we derive the near-zone 1.51.5PN dipole radiation–reaction force and combine it with the Darwin Hamiltonian through 11PN order to obtain a closed canonical $N$-body phase-space system. The equations are directly implementable, strictly conservative when dissipation is switched off, and exhibit monotonic energy loss, secular inspiral, and circularization when radiation reaction is included. For binaries we derive analytic circular and eccentric inspiral laws, including 11PN conservative corrections to the dipole-driven inspiral and merger time. Extending to charged compact binaries in Einstein–Maxwell theory, we combine the 22PN ADM-type conservative Hamiltonian with leading 1.51.5PN dipole dissipation and gravitational quadrupole flux, obtaining gauge-invariant energy–frequency relations, closed-form circular inspiral laws, and a dipole–quadrupole crossover scale that separates electromagnetic and gravitational flux dominated inspirals.
@@ I Introduction
Radiation reaction—the backreaction of emitted radiation on the motion of its source—is a unifying element of classical field theories, ranging from electrodynamics to the dissipative inspiral of compact binaries in general relativity [26, 22, 6, 16]. In gravitational-wave physics it is routinely incorporated within a canonical “Hamiltonian plus radiation–reaction” paradigm, implemented in PN/ADM and EOB frameworks [23, 24, 25, 27, 28, 29, 30, 31, 32] and complemented by nonconservative action/EFT formulations [33, 34]. In electrodynamics, a local self-force law arises from Dirac’s world-tube momentum balance and equivalent radiative-field decompositions, leading to the covariant Lorentz–Dirac (LD) equation [7, 9, 8, 10]. Its third-order character admits runaway and preaccelerated solutions, and the standard resolution is Landau–Lifshitz (LL) reduction of order, which yields a causal second-order dynamics perturbatively equivalent to LD [10, 13, 11, 12, 15]. These structural analogies motivate the main goal of this work: to construct an explicit, directly implementable electromagnetic analogue of the PN canonical paradigm and to use it for analytic and numerical studies for dissipative long-range dynamics.
We begin by adopting the order-reduced LL dynamics and taking its nonrelativistic near-zone limit, where dissipation first enters at absolute 1.51.5PN order and is governed by the electric dipole moment, yielding a dipole radiation–reaction force consistent with Larmor energy balance (up to the Schott term) [10, 8, 11]. We then build an $N$-body phase-space system in direct analogy with PN gravity: a conservative Hamiltonian sector truncated at Darwin (11PN) order plus an explicit non-Hamiltonian dipole radiation–reaction force at 1.51.5PN order, expressed entirely in canonical variables. This construction is designed to be operational (ready for numerical integration) and diagnostically transparent: in the conservative limit it preserves the Darwin Hamiltonian without secular drift, while with radiation reaction it exhibits monotonic energy loss, secular inspiral, and robust circularization. Specializing to $N=2$ we recover the characteristic dipole structure proportional to the charge-to-mass asymmetry and the exact suppression of the 1.51.5PN radiation reaction when $q_{1}/m_{1}=q_{2}/m_{2}$, in agreement with post-Coulombian analyses [35]. Using the same dissipative input, we derive analytic inspiral laws for circular and eccentric binaries, including secular evolution and closed-form integrals for $a(e)$ and $t(e)$ at leading order, and we validate these results against direct integrations of the full canonical equations.
To connect this purely electromagnetic system to relativistic charged compact binaries, we also examine inspiral laws in Einstein–Maxwell theory by combining the 22PN ADM-type center-of-mass Hamiltonian with leading 1.51.5PN dipole dissipation, following recent results for charged black-hole binaries [36]. This yields circular energy–frequency relations in terms of an effective coupling and analytic circular inspiral laws through 22PN conservative order. Including gravitational quadrupole flux in the balance law then identifies a dipole–quadrupole crossover scale that depends only on charge-to-mass asymmetry, interpolating between the dipole scaling $\dot{\Omega}\propto\Omega^{3}$ and the GR quadrupole scaling $\dot{\Omega}\propto\Omega^{11/3}$ [26]. We estimate the dipole–quadrupole crossover frequency as $f_{\rm cross}\simeq 2.2\times 10^{3}\,\mathrm{Hz}\,(M_{\odot}/M)\,|\eta_{2}-\eta_{1}|^{3}/(1-\eta_{1}\eta_{2})$, where $M$ is the total mass and $\eta_{A}\equiv\sqrt{k_{E}/G}\,(q_{A}/m_{A})$ denotes the dimensionless charge-to-mass ratio of body $A$. For a $60\,M_{\odot}$ binary this reduces to $f_{\rm cross}\simeq 36\,\mathrm{Hz}\,|\eta_{2}-\eta_{1}|^{3}/(1-\eta_{1}\eta_{2})$. The cubic dependence on the charge-to-mass asymmetry implies that for $|\eta_{A}|\ll 1$ the crossover lies far below the ground-based detector band, so the inspiral remains quadrupole dominated; dipole radiation effects enter the LIGO/Virgo band only if the charge-to-mass asymmetry $|\eta_{2}-\eta_{1}|$ is of order unity, i.e. in highly asymmetric and near-extremal configurations. It is also noteworthy that order-reduced electromagnetic radiation-reaction dynamics admits special analytic structures in certain limits, including exact reductions to Painlevé transcendents in Coulombic settings [37, 38]. Such features have motivated recent conjectures on universality in binary black-hole coalescence [39, 40, 41]. The present framework provides an explicit relativistic realization of dipole-driven inspiral within an explicit post-Newtonian structure.
@@
In Sec. II we review radiation reaction for point charges, deriving the Lorentz–Dirac equation, implementing the Landau–Lifshitz order-reduction procedure, and extracting the 1.51.5PN near-zone dipole radiation–reaction force. Section III combines this dissipative sector with the Darwin Hamiltonian to construct a closed 11PN+1.5+1.5PN canonical $N$-body phase-space system. In Sec. IV we specialize to $N=2$, derive the explicit dipole radiation–reaction structure for binaries, and connect with existing post-Coulombian and Einstein–Maxwell results. Sections V and VI develop analytic circular and eccentric inspiral laws at leading dipole order, while Sec. VII incorporates 11PN conservative corrections into the eccentric evolution. We then extend the framework to relativistic charged binaries in Einstein–Maxwell theory: Sec. VIII presents the ADM-type center-of-mass Hamiltonian through 22PN order together with the leading 1.51.5PN radiation reaction. Section IX analyzes the dipole–quadrupole crossover scale, and Secs. X and XI derive the circular inspiral laws through 22PN conservative order, first in the dipole-dominated regime and then including both dipole and quadrupole radiation.
Technical details and validation material are collected in the Appendices: Appendix A discusses runaway solutions of the Lorentz–Dirac equation; Appendix B presents single-particle consistency tests of the Landau–Lifshitz dynamics; Appendix C provides details of the 11PN Darwin force correction; and Appendix D summarizes numerical simulations of the charge-neutral binary.
II Radiation reaction: Lorentz–Dirac, Landau–Lifshitz reduction, and the near-zone limit
Radiation reaction for a point charge may be derived either by isolating the radiative field via the half–retarded minus half–advanced decomposition [10, 8] or by Dirac’s world–tube momentum balance [7, 9]. Both routes lead to the covariant Lorentz–Dirac (LD) equation
$ma^{\alpha}=F^{\alpha}_{\rm ext}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta},$ (1)
where $u^{\alpha}$ is the four–velocity, $a^{\alpha}\!=\!du^{\alpha}/d\tau$, and the projector enforces $u_{\alpha}a^{\alpha}=0$ (with $u_{\alpha}u^{\alpha}=-c^{2}$). Equation (1) is third order and admits runaway and preaccelerated solutions; this is already evident in the nonrelativistic limit,
$a(t)-t_{0}\,\dot{a}(t)=\frac{1}{m}F_{\rm ext}(t),\qquad t_{0}=\frac{1}{4\pi\varepsilon_{0}}\,\frac{2q^{2}}{3mc^{3}},$ (2)
for which a step force $F_{\rm ext}(t)=f\,\theta(t)$ yields
$a(t)=e^{t/t_{0}}\!\left[b-\frac{f}{m}\bigl(1-e^{-t/t_{0}}\bigr)\theta(t)\right],\qquad a(t)=\frac{f}{m}\left[\theta(-t)e^{t/t_{0}}+\theta(t)\right]\ \ (b=f/m),$ (3)
showing, respectively, runaway growth for generic $b$ and preacceleration for the tuned (no–runaway) choice.
The physical dynamics is obtained by the Landau–Lifshitz (LL) reduction of order [10, 13], treating the self–force perturbatively and replacing $\dot{a}^{\alpha}$ by the proper–time derivative of the leading Lorentz–force acceleration. This gives the second–order LL equation
@@ II Radiation reaction: Lorentz–Dirac, Landau–Lifshitz reduction, and the near-zone limit
$ma^{\alpha}=F^{\alpha}_{\rm ext}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}mc^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)F^{\beta}_{{\rm ext},\gamma}u^{\gamma},$ (4)
which is causal and free of runaway/preacceleration pathologies [13, 12, 11]. For electromagnetic forcing $F^{\alpha}_{\rm ext}=q\,F^{\alpha}{}_{{\rm ext}\,\mu}u^{\mu}$, one convenient explicit form is
$ma^{\alpha}$ $=q\,F^{\alpha}{}_{{\rm ext}\,\mu}\,u^{\mu}+q\,t_{0}\,F^{\alpha}{}_{{\rm ext}\,\mu,\nu}\,u^{\mu}u^{\nu}+\frac{q^{2}t_{0}}{m}\left(\delta^{\alpha}{}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)F^{\beta}{}_{{\rm ext}\,\mu}\,F^{\mu}{}_{{\rm ext}\,\nu}\,u^{\nu},$ (5)
with $t_{0}$ given in (2) (one intermediate term vanishes by antisymmetry of $F_{\mu\nu}$). In Dirac’s world–tube picture, the electromagnetic momentum flux through a tube $\Sigma$ of radius $r$,
$\frac{dP^{\alpha}_{\rm em}}{d\tau}=\int_{\Sigma}T^{\alpha\beta}_{\rm em}\,d\Sigma_{\beta},$ (6)
splits into a divergent near–field piece and a finite radiative piece,
$\frac{dP^{\alpha}_{\rm em}}{d\tau}=\frac{q^{2}}{8\pi\varepsilon_{0}rc^{2}}\,a^{\alpha}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\frac{a^{2}}{c^{2}}\,u^{\alpha},$ (7)
where the $1/r$ term is absorbed into the renormalized mass and the finite term encodes radiated energy–momentum. The equivalent “Dirac” form of the LD equation,
$ma^{\alpha}=F^{\alpha}_{\rm ext}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\left(\dot{a}^{\alpha}-\frac{a^{2}}{c^{2}}u^{\alpha}\right),$ (8)
reduces to (1) upon using